OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the point (6, 2) and has a slope of 3?

OpenStudy (anonymous):

Help please :)

OpenStudy (anonymous):

@Destinymasha

OpenStudy (anonymous):

@undeadknight26 @rose21 @Mimi_x3 @pooja195 @Compassionate @chosenmatt @calculusxy @Jesstho.-. @micahm @KamiBug @dan815 @dan815 @tHe_FiZiCx99 @TheSmartOne

OpenStudy (micahm):

en.wikipedia.org/wiki/Linear_equation Wikipedia Linear equations can have one or more variables. ... 2.1.2 Slope–intercept form; 2.1.3 Point–slope form; 2.1.4 Two-point form ... Since terms of linear equations cannot contain products of distinct or equal variables, nor any power (other than 1) or other function of a variable, equations involving terms such as xy, x2, y1/3, and ...

OpenStudy (anonymous):

Do you have to tag this many people?

OpenStudy (anonymous):

^ yes :/

OpenStudy (anonymous):

will fan for answrrr

OpenStudy (anonymous):

@HelpBlahBlahBlah @CausticSyndicalist @mrdoldum @Goku-Kai @Gin_Ichimaru @LiteNing1337 @jswagger @Here_to_Help15 @bohotness

OpenStudy (anonymous):

y = mx + b 1 = 6(1) + b 1 = 6 + (-5) y = 6x - 5

OpenStudy (anonymous):

Did you change it to wumbo?

OpenStudy (anonymous):

thank you jswagger

OpenStudy (bohotness):

huh

OpenStudy (anonymous):

welcome boo